### v1

# The late Miocene 10Be anomaly and the possibility of a supernova

11email: efrem.maconi@univie.ac.at 22institutetext: Astronomical Institute of the Czech Academy of Sciences, Boční II 1401, 141 31 Prague 4, Czech Republic 33institutetext: Center for Astrophysics — Harvard & Smithsonian, 60 Garden St., Cambridge, MA 02138, USA
The late Miocene 10Be anomaly and the possibility of a supernova
Recent measurements of cosmogenic 10Be in deep-ocean ferromanganese crusts from the Central and Northern Pacific have revealed an anomalous concentration between 11.5 and 9.0 Myr ago, peaking at 10.1 Myr. One possible explanation is a nearby supernova (SN) event. Motivated by this and by the proximity of the Solar System to the Orion star-forming region during that period, we estimate the probability that at least one SN occurred between the onset and peak of the anomaly. Using an open cluster catalog based on Gaia DR3, we trace back the orbits of 2725 clusters and the Sun over the past 20 Myr and compute the expected number of SN events. We find 19 clusters with a probability greater than 1% each of producing at least one SN within 100 pc of the Sun in the time interval 11.5–10.1 Myr ago. The total cumulative probability exceeds zero at 35 pc from the Sun and increases rapidly with distance, reaching 68% near 100 pc. Two young clusters dominate the SN probability: ASCC 20 contributes most within 70 pc, while OCSN 61 becomes more significant beyond that distance. Our results support the plausibility of a SN origin for the 10Be anomaly and highlight the importance of additional 10Be records from independent terrestrial archives to determine whether the anomaly is of astrophysical or terrestrial origin.
Key Words.:
Galaxy: solar neighborhood - ISM: kinematics and dynamics - open clusters and associations: general1 Introduction
The Solar System orbits the center of the Milky Way, together with billions of other stars and vast reservoirs of interstellar gas. Characterizing past environments crossed by the Sun, including possible encounters with large-scale Galactic structures or nearby supernovae (SNe), helps to find possible connections between Galactic environments and Earth’s geological record, fostering interdisciplinary research (see e.g., Fuchs et al., 2006; Breitschwerdt et al., 2016; Koll et al., 2019; Wallner et al., 2021; Miller & Fields, 2022; Opher et al., 2024; Maconi et al., 2025; Zucker et al., 2025).
From an astronomical perspective, the European Space Agency’s Gaia mission (Gaia Collaboration et al., 2016) has enabled significant advancements, revolutionizing our understanding of the local Galactic environment. For example, Gaia data have been used to unveil the 3D structure of the solar neighborhood (see e.g., Leike et al., 2020; Vergely et al., 2022; Edenhofer et al., 2024), compile new molecular cloud catalogs (see e.g., Zucker et al., 2019; Cahlon et al., 2024), and significantly expand the open cluster census (see e.g., Castro-Ginard et al., 2018; Cantat-Gaudin et al., 2018; Hunt & Reffert, 2023). These achievements have provided new insights into the formation and evolution of stellar clusters (see e.g., Meingast & Alves, 2019; Swiggum et al., 2024), helped constrain the structure and history of the Local Bubble (Zucker et al., 2022; O’Neill et al., 2024), and led to the identification of previously unknown Galactic structures, such as the Radcliffe wave (see e.g., Alves et al., 2020), along with their relation to the past trajectory of the Solar System (see e.g., Maconi et al., 2025).
Concurrent to these astronomical advancements, studies of long-lived radionuclides, like 60Fe (; Wallner et al. 2015a) in geological archives have revealed signatures indicative of nearby SNe or encounters with interstellar regions enriched with these elements (see e.g., Knie et al., 1999; Wallner et al., 2015b, 2021; Koll et al., 2019). In addition, cosmogenic nuclides such as 14C (; Godwin 1962) and 10Be (; Chmeleff et al. 2010), are used for archaeological and geological dating on kyr to Myr timescales. Anomalies in their concentration profiles are significant not only for their physical interpretation but also because, if found in multiple independent archives, they may serve as chronological anchor markers (see e.g., Dee & Pope, 2016).
Recently, Koll et al. (2025) reported on the discovery of a 10Be anomaly in deep ocean crusts of the Central and Northern Pacific during the late Miocene. The origin of this anomaly, dated between 11.5 and 9.0 Myr ago and peaking at 10.1 Myr, remains uncertain, and several scenarios are discussed by the authors. As 10Be is produced by cosmic-ray (CR) spallation in the upper atmosphere (see e.g., Webber & Higbie, 2003), one possibility is that a nearby SN may be responsible for the 10Be excess.
In a recent study, Maconi et al. (2025) showed that around 11.5 Myr ago, at the onset of the 10Be anomaly, the Solar System was exiting the Radcliffe Wave, leaving behind the Orion star formation region, a major site of recent star formation. Given the Sun’s proximity at that time to several massive young clusters, a nearby SN event is a plausible explanation for the observed 10Be anomaly. In this work, we test this hypothesis by integrating the orbits of the Sun and a large sample of clusters back in time over the past 20 Myr and by estimating the probability that a SN occurred within a given distance of the Solar System during the 10Be anomaly.
2 Data
We primarily used the open cluster catalog by Hunt & Reffert (2023), which was constructed by using the HDBSCAN algorithm (McInnes et al., 2017) on the Gaia DR3 astrometric data (Gaia Collaboration et al., 2023). The catalog contains a total of 7166 star clusters and provides a wide range of parameters, including sky positions, parallaxes, proper motions, extinctions, and stellar membership lists (see Hunt & Reffert, 2023, for details).
We complemented this catalog with four additional clusters not included in Hunt & Reffert (2023), namely CWNU 1028, NGC 1977, OC 0340, and UBC 207. These clusters are part of the Radcliffe Wave and were identified and used in previous studies (Konietzka et al., 2024; Maconi et al., 2025), as detailed in Appendix A. Moreover, we updated the stellar membership of six clusters in the Orion region (ASCC 19, ASCC 20, OCSN 56, OCSN 61, OCSN 65, and Theia 13) with additional members identified using the Significance Mode Analysis (SigMA) clustering algorithm (see Ratzenböck et al., 2023a), applied specifically to the Orion region (A. Rottensteiner, in preparation).
By crossmatching individual cluster members with supplementary radial velocity (RV) surveys, we were able to improve the accuracy of the mean kinematic data of the clusters over the Gaia-only values. For our purposes, we only considered clusters with reliable RV measurements () available for at least three member stars per cluster. After imposing these kinematic quality criteria, our sample comprises of 2725 clusters. Further details on the data curation process are provided in Appendix A.
3 Methods
3.1 Selection of the clusters of interest
We performed a preliminary orbital integration (see Sect. 3.3) of our initial cluster sample, identifying those that approached the Solar System within a threshold distance of 200 pc over the past 20 Myr. This threshold was chosen to ensure completeness, considering that SNe occurring within 100-150 pc may leave detectable traces on Earth (see e.g., Fry et al., 2015). This step reduced the sample to 278 clusters, about 10% of the initial sample.
For each of these 278 clusters, we performed 1000 orbital integrations, varying their initial positions and velocities by Monte Carlo (MC) sampling the uncertainty distributions. We then selected only the clusters that approached the Solar System within 100 pc during the time span of the 10Be anomaly (9–11.5 Myr ago; see Koll et al., 2025). We adopt 100 pc as a conservative threshold, balancing the range at which SNe may leave traces on Earth with the need for proximity to explain the 10Be anomaly (Koll et al., 2025). We do not explicitly account for the travel time between the SN event and the arrival of CRs at Earth, as this interval is negligible compared to the Myr-scale intervals studied here. This resulted in a subset of 71 clusters.
3.2 Age, mass, and SNe estimation
For the subset of 71 selected clusters identified in Sect. 3.1, we estimated ages, present-day masses, completeness-corrected masses, and the probability of SN events. For the age estimation, we used the Chronos111Chronos code: https://github.com/sebastianratzenboeck/Chronos Python package (see Ratzenböck et al., 2023b), which performs a Bayesian fit of theoretical isochrones to the color magnitude diagram of each cluster. We used the PARSEC isochrones models (Bressan et al., 2012; Nguyen et al., 2022) and assumed solar metallicity () for the stars in the clusters. We limited our fit to to as stars fainter than this threshold have been empirically found to be less well described by isochrone models (see e.g., Ratzenböck et al., 2023b; Rottensteiner & Meingast, 2024).
The present-day cluster masses were estimated by adding up the masses of its stellar members, derived from the corresponding best-fit isochrone. We also computed the initial mass of the clusters corrected for incompleteness, as some star members may be missing due to observational limits, limitation of the clustering algorithm, and stellar evolution (see e.g., Meingast et al., 2021; Hunt & Reffert, 2023; Ratzenböck et al., 2023a): following Meingast et al. (2021), we minimized the difference between the present-day mass function and a Kroupa (2001) initial mass function (IMF). To account for the completeness limits of Gaia data (Gaia Collaboration et al., 2023), we performed the minimization within the mass range . The low-mass end limit of the Kroupa IMF was set to to account for objects below the hydrogen-burning limit, while no high-mass bound was imposed to account for massive stars potentially absent due to Gaia brightness limit or past SNe. The IMF was sampled using the IMF Python code 222IMF code: https://github.com/keflavich/imf.
For each cluster, we then estimated the number of SNe that may have occurred while it was within 100 pc of the Solar System and during the time interval between the onset and the peak of the 10Be anomaly (11.5-10.1 Myr ago). This corresponds to the period during which a nearby SN would be most relevant to explain the peak in 10Be. For each of the 1000 orbital realizations (see Sect. 3.3), we assigned an age and a mass to the cluster by sampling their respective posterior distributions as provided by Chronos. The sampled mass was used to generate a synthetic stellar population assuming a Kroupa IMF. The sampled age was used to compute the age of the cluster at the beginning and end of its proximity to the Sun within the time window of interest. We then counted the number of massive stars () with masses greater than the most massive star predicted by the PARSEC stellar evolutionary model for the corresponding ages. From this, we derived the probability of hosting at least one SN event during the considered time range and across different distance thresholds, both for each cluster individually and for the entire cluster ensemble. This procedure was repeated 100 times to estimate the statistical uncertainties on the probabilities.
3.3 Integration of the orbits
We computed the past orbital trajectories of the clusters and the Solar System using the Galactic dynamics package galpy (Bovy, 2015). galpy enables the numerical integration of orbits over different initial conditions (i.e., Galactocentric distance, height of the Sun above the Galactic disk, velocity of the Sun, and velocity of the Local Standard of Rest; LSR) and various models for the Milky Way potential.
For this study, we adopted the MWPotential2014 model offered by galpy as the gravitational potential of the Milky Way. This model includes a bulge, disk, and dark-matter halo component (see Bovy, 2015, for details). We assumed a solar Galactocentric distance of (Gillessen et al., 2009) and a solar height of (Chen et al., 2001). The velocity of the LSR was set to and the velocity of the Sun relative to it to () = (11.1, 12.24, 7.25) (Schönrich et al., 2010). Orbital integrations were performed over the past 20 Myr with a 0.01 Myr time step using the dop853-c Dormand–Prince integrator in galpy for computational efficiency.
Statistical uncertainties in the positions and velocities of the clusters and of the Sun were addressed by repeating the orbital integration procedure 1000 times, each time using a new realization of the input data by MC sampling the associated uncertainties. In Appendix C, we assess the impact of adopting a different set of Solar parameters and find that the relative distances of the Sun and clusters remain largely unchanged over the past 20 Myr, supporting the robustness of our conclusions.
4 Results and Discussion
Various scenarios are considered by Koll et al. (2025) to explain the 10Be anomaly in the late Miocene. Some involve geological processes that could increase 10Be concentrations in ocean water without altering its atmospheric production rate, while others explore the possibility of an actual enhancement in Galactic CR flux reaching Earth. Among the most promising explanations are the onset and intensification of the Antarctic Circumpolar Current as a terrestrial cause, and either a nearby SN or the compression of the heliosphere by the passage of the Sun through a dense interstellar cloud as astrophysical origins. We refer the reader to their work for a more detailed discussion of all proposed scenarios and focus here solely on the SN hypothesis, further motivated by the fact that around the onset of the 10Be anomaly (11.5 Myr ago), the Solar System was leaving the Orion region of the Radcliffe Wave behind (Maconi et al., 2025), and was therefore in proximity to an active star-forming region where several massive clusters were forming or had just formed.
We find that, out of the 2725 open clusters in our initial sample (see Sect. 2), 19 have a probability greater than 1% of hosting at least one SN within 100 pc of the Solar System and between the onset and the peak of the 10Be anomaly (11.5–10.1 Myr ago), and only four exceed this threshold within 70 pc, namely ASCC 20, OCSN 61, OCSN 56, and CWNU 1057. We estimate the total probability of at least one SN event within 35 pc of the Sun to be around 1%, rising to 5.4% at 40 pc, 14% at 50 pc, and 28% at 70 pc. At 100 pc the total probability across all clusters reaches 68%. We note that none of the considered clusters come closer than 20 pc, thus not reaching the critical distance (8–20 pc) within which a SN could cause an extinction event (see e.g., Gehrels et al., 2003; Thomas & Yelland, 2023). In Fig. 1, we show the distances between the four mentioned clusters and the Solar System over the past 20 Myr, their individual SN probabilities, the total SN probability from all clusters during the relevant interval, and the 10Be concentration profile by Koll et al. (2025). In Table 3, we list the 19 clusters of interest together with their SN probability for different threshold distances. Their properties, including ages and masses, are reported in Table 2, while their heliocentric positions and velocities are given in Table 1.
Among the 19 clusters, ASCC 20 and OCSN 61 (also known as OBP-b), both located in Orion, are the dominant contributors. ASCC 20 is the only cluster with a significant contribution within 70 pc (up to %), while OCSN 61 becomes increasingly relevant beyond 70 pc, reaching 29% at 100 pc. For ASCC 20, which reaches a minimum distance of from the Sun around 11.8 Myr ago and remains within 100 pc throughout the 10Be anomaly, we estimated an age of , consistent with values reported in the literature (see e.g., Kos et al., 2019; Maconi et al., 2025), and derive a present day mass of and an incompleteness corrected initial mass of . For OCSN 61, which remains within 100 pc between the onset and peak of the anomaly but never approaches the Sun closer than 60 pc, we estimated an age of , a present day mass of , and an incompleteness corrected initial mass of . For both of these clusters, we updated the stellar membership using the SigMA algorithm (see Sect. 2 and Appendix A), identifying new members compared to the one listed in Hunt & Reffert (2023). We note that even with an improved member selection, the cluster catalogs are still likely incomplete, suggesting that the estimated number of SN events is conservative due to the possible incompleteness of the cluster catalogs.
To account for systematic uncertainties in the geological age dating, we adopt a conservative estimate of Myr, based on the largest uncertainty reported for one of the samples (Crust-3) used by Koll et al. (2025) and previously analyzed by Wallner et al. (2021). We repeated our analysis with two shifted time windows: [12.0, 10.6] Myr ago and [11.0, 9.6] Myr ago. We find that the total SN probability slightly increases for the older time window and slightly decreases for the younger one, driven primarily by the changing in proximity of ASCC 20 and OCSN 61 (see Fig. 2 and Appendix D). These variations do not affect the result that a nearby SN remains a plausible explanation for the 10Be anomaly.
We note that the 10Be anomaly appears as a broad peak rather than a sharp one, as might be expected from a SN. However, the time interval derived from the crust might be broader than the actual signal, due to diffusion and redistribution processes. Moreover, Koll et al. (2025) consider the SN hypothesis a viable scenario only under certain energetic and geometrical conditions. We consider this plausible given the many uncertainties and physical processes involved in a SN and CR production, including the total SN energy (see e.g., Kasen & Woosley, 2009), the efficiency of energy conversion into CR (see e.g., Blasi, 2011), as well as CR acceleration (see e.g., Caprioli, 2012) and transport (see e.g., Amato & Blasi, 2018). Future advancements in the understanding of these processes will help to better constrain the conditions under which a SN could produce such an anomaly.
The presence of other radionuclides could offer further insights into the origin of the 10Be anomaly and potentially support the nearby SN scenario. For example, 60Fe, produced in SNe, has been detected in geological archives and linked to past SN events (Wallner et al., 2021). So far, there is no indication of a concomitant 60Fe signal with the 10Be anomaly. However, 10Be enhancements have not been observed in association with previously reported 60Fe peaks and this apparent lack of correlation remains an open question. Additionally, the detection of 60Fe within the time window considered here is further complicated by its advanced decay and the expected low concentrations. Among other cosmogenic radionuclides, 53Mn () is a valuable candidate for complementary investigation alongside 10Be, and future accelerator mass spectrometry facilities may enable its detection in the relevant time window (see Koll et al., 2025, for further discussion).
In conclusion, we find that a nearby SN remains a plausible explanation for the 10Be anomaly, especially given the Solar System’s proximity to the Orion region during that period. The estimated SN probability is nonzero at 35 pc and increases with distance, with ASCC 20 and OCSN 61 emerging as the most promising candidate clusters. ASCC 20 is the primary contributor up to 70 pc, while OCSN 61 becomes more relevant beyond that distance. Future investigations of 10Be records from terrestrial archives outside the Pacific Ocean will be crucial to determine whether the observed anomaly reflects a global signal or a regional effect confined to this basin, helping to constrain its terrestrial or astrophysical origin.
Acknowledgements.
EM, JA, and AR were co-funded by the European Union (ERC, ISM-FLOW, 101055318). JG was co-funded from the European Union, Central Bohemian Region, and Czech Academy of Sciences, as part of the MERIT fellowship (MSCA-COFUND Horizon Europe, Grant agreement 101081195). This work has made use of data from the European Space Agency (ESA) mission Gaia https://www.cosmos.esa.int/gaia, processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.References
- Abdurro’uf et al. (2022) Abdurro’uf, Accetta, K., Aerts, C., et al. 2022, ApJS, 259, 35
- Alves et al. (2020) Alves, J., Zucker, C., Goodman, A. A., et al. 2020, Nature, 578, 237
- Amato & Blasi (2018) Amato, E. & Blasi, P. 2018, Advances in Space Research, 62, 2731
- Bennett & Bovy (2019) Bennett, M. & Bovy, J. 2019, MNRAS, 482, 1417
- Blasi (2011) Blasi, P. 2011, in Cosmic Rays for Particle and Astroparticle Physics, ed. S. Giani, C. Leroy, & P. G. Rancoita, 493–506
- Bovy (2015) Bovy, J. 2015, ApJS, 216, 29
- Breitschwerdt et al. (2016) Breitschwerdt, D., Feige, J., Schulreich, M. M., et al. 2016, Nature, 532, 73
- Bressan et al. (2012) Bressan, A., Marigo, P., Girardi, L., et al. 2012, MNRAS, 427, 127
- Buder et al. (2025) Buder, S., Kos, J., Wang, X. E., et al. 2025, PASA, 42, e051
- Cahlon et al. (2024) Cahlon, S., Zucker, C., Goodman, A., Lada, C., & Alves, J. 2024, ApJ, 961, 153
- Cantat-Gaudin et al. (2020) Cantat-Gaudin, T., Anders, F., Castro-Ginard, A., et al. 2020, A&A, 640, A1
- Cantat-Gaudin et al. (2018) Cantat-Gaudin, T., Vallenari, A., Sordo, R., et al. 2018, A&A, 615, A49
- Caprioli (2012) Caprioli, D. 2012, J. Cosmology Astropart. Phys., 2012, 038
- Castro-Ginard et al. (2018) Castro-Ginard, A., Jordi, C., Luri, X., et al. 2018, A&A, 618, A59
- Chen et al. (2001) Chen, B., Stoughton, C., Smith, J. A., et al. 2001, ApJ, 553, 184
- Chmeleff et al. (2010) Chmeleff, J., von Blanckenburg, F., Kossert, K., & Jakob, D. 2010, Nuclear Instruments and Methods in Physics Research B, 268, 192
- Dee & Pope (2016) Dee, M. W. & Pope, B. J. S. 2016, Proceedings of the Royal Society of London Series A, 472, 20160263
- Edenhofer et al. (2024) Edenhofer, G., Zucker, C., Frank, P., et al. 2024, A&A, 685, A82
- Fry et al. (2015) Fry, B. J., Fields, B. D., & Ellis, J. R. 2015, ApJ, 800, 71
- Fuchs et al. (2006) Fuchs, B., Breitschwerdt, D., de Avillez, M. A., Dettbarn, C., & Flynn, C. 2006, MNRAS, 373, 993
- Gaia Collaboration et al. (2016) Gaia Collaboration, Prusti, T., de Bruijne, J. H. J., et al. 2016, A&A, 595, A1
- Gaia Collaboration et al. (2023) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023, A&A, 674, A1
- Gehrels et al. (2003) Gehrels, N., Laird, C. M., Jackman, C. H., et al. 2003, ApJ, 585, 1169
- Gillessen et al. (2009) Gillessen, S., Eisenhauer, F., Trippe, S., et al. 2009, ApJ, 692, 1075
- Godwin (1962) Godwin, H. 1962, Nature, 195, 984
- Gontcharov (2006) Gontcharov, G. A. 2006, Astronomical and Astrophysical Transactions, 25, 145
- GRAVITY Collaboration et al. (2018) GRAVITY Collaboration, Abuter, R., Amorim, A., et al. 2018, A&A, 615, L15
- Hao et al. (2022) Hao, C. J., Xu, Y., Wu, Z. Y., et al. 2022, A&A, 660, A4
- He et al. (2022) He, Z., Wang, K., Luo, Y., et al. 2022, ApJS, 262, 7
- Hunt & Reffert (2023) Hunt, E. L. & Reffert, S. 2023, A&A, 673, A114
- Kasen & Woosley (2009) Kasen, D. & Woosley, S. E. 2009, ApJ, 703, 2205
- Knie et al. (1999) Knie, K., Korschinek, G., Faestermann, T., et al. 1999, Phys. Rev. Lett., 83, 18
- Koll et al. (2019) Koll, D., Korschinek, G., Faestermann, T., et al. 2019, Phys. Rev. Lett., 123, 072701
- Koll et al. (2025) Koll, D., Lachner, J., Beutner, S., et al. 2025, Nature Communications, 16, 866
- Konietzka et al. (2024) Konietzka, R., Goodman, A. A., Zucker, C., et al. 2024, Nature, 628, 62
- Kos et al. (2019) Kos, J., Bland-Hawthorn, J., Asplund, M., et al. 2019, A&A, 631, A166
- Kroupa (2001) Kroupa, P. 2001, MNRAS, 322, 231
- Leike et al. (2020) Leike, R. H., Glatzle, M., & Enßlin, T. A. 2020, A&A, 639, A138
- Liu & Pang (2019) Liu, L. & Pang, X. 2019, ApJS, 245, 32
- Maconi et al. (2025) Maconi, E., Alves, J., Swiggum, C., et al. 2025, A&A, 694, A167
- McInnes et al. (2017) McInnes, L., Healy, J., & Astels, S. 2017, The Journal of Open Source Software, 2, 205
- Meingast & Alves (2019) Meingast, S. & Alves, J. 2019, A&A, 621, L3
- Meingast et al. (2021) Meingast, S., Alves, J., & Rottensteiner, A. 2021, A&A, 645, A84
- Miller & Fields (2022) Miller, J. A. & Fields, B. D. 2022, ApJ, 934, 32
- Miret-Roig et al. (2020) Miret-Roig, N., Galli, P. A. B., Brandner, W., et al. 2020, A&A, 642, A179
- Nguyen et al. (2022) Nguyen, C. T., Costa, G., Girardi, L., et al. 2022, A&A, 665, A126
- O’Neill et al. (2024) O’Neill, T. J., Zucker, C., Goodman, A. A., & Edenhofer, G. 2024, arXiv e-prints, arXiv:2403.04961
- Opher et al. (2024) Opher, M., Loeb, A., & Peek, J. E. G. 2024, Nature Astronomy
- Randich et al. (2022) Randich, S., Gilmore, G., Magrini, L., et al. 2022, A&A, 666, A121
- Ratzenböck et al. (2023b) Ratzenböck, S., Großschedl, J. E., Alves, J., et al. 2023b, A&A, 678, A71
- Ratzenböck et al. (2023a) Ratzenböck, S., Großschedl, J. E., Möller, T., et al. 2023a, A&A, 677, A59
- Rottensteiner & Meingast (2024) Rottensteiner, A. & Meingast, S. 2024, A&A, 690, A16
- Schönrich et al. (2010) Schönrich, R., Binney, J., & Dehnen, W. 2010, MNRAS, 403, 1829
- Sim et al. (2019) Sim, G., Lee, S. H., Ann, H. B., & Kim, S. 2019, Journal of Korean Astronomical Society, 52, 145
- Steinmetz et al. (2020) Steinmetz, M., Matijevič, G., Enke, H., et al. 2020, AJ, 160, 82
- Swiggum et al. (2024) Swiggum, C., Alves, J., Benjamin, R., et al. 2024, Nature, 631, 49
- Szilágyi et al. (2021) Szilágyi, M., Kun, M., & Ábrahám, P. 2021, MNRAS, 505, 5164
- Thomas & Yelland (2023) Thomas, B. C. & Yelland, A. M. 2023, ApJ, 950, 41
- Torres et al. (2006) Torres, C. A. O., Quast, G. R., da Silva, L., et al. 2006, A&A, 460, 695
- Tsantaki et al. (2022) Tsantaki, M., Pancino, E., Marrese, P., et al. 2022, A&A, 659, A95
- Vergely et al. (2022) Vergely, J. L., Lallement, R., & Cox, N. L. J. 2022, A&A, 664, A174
- Wallner et al. (2015a) Wallner, A., Bichler, M., Buczak, K., et al. 2015a, Phys. Rev. Lett., 114, 041101
- Wallner et al. (2015b) Wallner, A., Faestermann, T., Feige, J., et al. 2015b, Nature Communications, 6, 5956
- Wallner et al. (2021) Wallner, A., Froehlich, M. B., Hotchkis, M. A. C., et al. 2021, Science, 372, 742
- Webber & Higbie (2003) Webber, W. R. & Higbie, P. R. 2003, Journal of Geophysical Research (Space Physics), 108, 1355
- Zhao et al. (2012) Zhao, G., Zhao, Y.-H., Chu, Y.-Q., Jing, Y.-P., & Deng, L.-C. 2012, Research in Astronomy and Astrophysics, 12, 723
- Zucker et al. (2022) Zucker, C., Goodman, A. A., Alves, J., et al. 2022, Nature, 601, 334
- Zucker et al. (2025) Zucker, C., Redfield, S., Starecheski, S., Konietzka, R., & Linsky, J. L. 2025, arXiv e-prints, arXiv:2504.00093
- Zucker et al. (2019) Zucker, C., Speagle, J. S., Schlafly, E. F., et al. 2019, ApJ, 879, 125
Appendix A Cluster catalog data
In this section, we describe the compilation of open clusters used in this study and explain how we refined the positional and kinematic data obtained from the primary catalog. As outlined in Sect. 2, we primary used the open cluster catalog by Hunt & Reffert (2023), which currently represents the largest open cluster search conducted in a homogeneous way. This catalog was constructed using the Hierarchical Density-Based Spatial Clustering of Applications with Noise routine (HDBSCAN; McInnes et al. 2017) on Gaia DR3 astrometric data (Gaia Collaboration et al. 2023), combined with a statistical density test and a Bayesian convolutional neural network for result validation (see Hunt & Reffert 2023, for details). The catalog contains a total of 7166 star clusters. Following Maconi et al. (2025), we added four additional clusters: CWNU 1028, NGC 1977, OC 0340, and UBC 207. These clusters were identified within other cluster catalogs (Sim et al. 2019; Liu & Pang 2019; Cantat-Gaudin et al. 2020; Szilágyi et al. 2021; Hao et al. 2022; He et al. 2022) which predated Hunt & Reffert (2023) and are considered to be part of the Radcliffe Wave (Konietzka et al. 2024; Maconi et al. 2025).
For six clusters in the Orion region (ASCC 19, ASCC 20, OCSN 56, OCSN 61, OCSN 65, and Theia 13), we were able to update the stellar membership using the SigMA clustering algorithm (Significance Mode Analysis; Ratzenböck et al. 2023a), applied to Gaia data and specifically tuned for the Orion complex (A. Rottensteiner, in preparation). The updated stellar membership for these six Orion clusters, compared to Hunt & Reffert (2023), changes as follows: ASCC 19 from 61 to 796, ASCC 20 from 194 to 525, OCSN 61 from 147 to 530, OCSN 65 from 70 to 534, OCSN 56 from 88 to 52, and Theia 13 from 249 to 221. The latter two have slightly fewer stellar members compared to Hunt & Reffert (2023). This can be explained by the fact that Hunt & Reffert (2023) prioritized precision over completeness, aiming to minimize the number of false positives associated with each cluster and to build a uniform catalog across the entire Solar neighborhood. In contrast, the updated stellar memberships were derived by applying SigMA specifically to the Orion region. We refer to Ratzenböck et al. (2023a), for a discussion on systematic biases coming from various clustering methods applied to the same region. The initial cluster sample comprises a total of 7170 members.
To obtain more accurate 3D space motions for each cluster, we supplement the Gaia DR3 RVs with additional RV data from supplementary surveys, including APOGEE-2 DR17 (Abdurro’uf et al. 2022), GALAH DR4 (Buder et al. 2025), RAVE DR6 (Steinmetz et al. 2020), Gaia ESO DR6 (Randich et al. 2022), two RV compilations (Gontcharov 2006; Torres et al. 2006), and LAMOST DR5 (Zhao et al. 2012, as corrected by Tsantaki et al. 2022). In cases where a star has multiple RV measurements from different surveys, we selected the RV value with the lowest uncertainty. Additionally, we only include sources with RV errors smaller than . To remove outliers, we applied 3-sigma-clipping around each clusters median RV value. We then derived the Cartesian velocities relative to the Sun () using the sub-sample of stars with valid RV measurements. For our study, we used medians in and , along with the associated uncertainties, as an estimate for the clusters bulk position and motion. Only clusters with at least three stars with RV measurements are included in our final sample, reducing the initial sample from 7170 to 2725.
Appendix B Cluster properties
In this Section we report the Tables that summarize the properties of the clusters as derived in the main part of the paper. The tables regards the 19 clusters that have a probability greater than 1% of hosting a SN within 100 pc of the Solar System and during the time interval 11.5–10.1 Myr ago, as described in Sects. 3 and 4.
Table 1 lists the Heliocentric Galactic Cartesian positions and velocities. Table 2 provides various information for the clusters, including ages, masses, minimum Sun–cluster distances, and time intervals during which they remain within 100 pc. Table 3 reports the SN probabilities for different threshold distances between 100 pc and 30 pc.
Name | ||||||||||||
ASCC 20 | -319.65 | -129.48 | -108.33 | -27.03 | -8.87 | -9.34 | 0.98 | 0.42 | 0.43 | 0.27 | 0.12 | 0.10 |
ASCC 24 | -164.72 | -124.71 | -26.95 | -13.65 | -9.72 | -10.00 | 0.93 | 0.97 | 0.83 | 0.61 | 0.39 | 0.19 |
CWNU 1057 | -114.40 | -63.81 | -54.58 | -13.45 | -9.39 | -7.95 | 0.62 | 1.38 | 0.56 | 1.10 | 0.61 | 0.57 |
CWNU 1111 | -256.25 | -56.66 | -42.29 | -19.22 | -9.50 | -9.33 | 1.74 | 1.03 | 0.41 | 1.67 | 0.44 | 0.25 |
HSC 1340 | -112.20 | 5.24 | -39.43 | -13.89 | -6.32 | -9.90 | 0.56 | 0.83 | 0.50 | 0.53 | 0.08 | 0.20 |
HSC 1373 | -107.87 | 29.02 | -82.49 | -12.43 | -5.41 | -5.95 | 1.46 | 2.50 | 0.83 | 0.91 | 0.29 | 0.99 |
HSC 1469 | -122.45 | -17.54 | -128.58 | -4.72 | -6.11 | -7.85 | 1.58 | 1.25 | 0.52 | 0.84 | 0.16 | 0.94 |
HSC 1523 | -257.27 | -56.94 | -75.51 | -16.28 | -9.18 | -6.13 | 1.68 | 1.03 | 0.66 | 2.04 | 0.43 | 0.60 |
HSC 1640 | -170.42 | -85.42 | -135.24 | -6.83 | -8.63 | -6.01 | 2.29 | 1.39 | 1.15 | 0.88 | 0.36 | 0.56 |
HSC 1687 | -312.44 | -208.38 | -46.00 | -23.38 | -24.10 | -6.82 | 1.48 | 2.16 | 0.74 | 0.18 | 0.16 | 0.18 |
HSC 1692 | -197.27 | -137.93 | -101.25 | -22.99 | -5.87 | -6.85 | 3.75 | 2.58 | 1.65 | 0.65 | 0.30 | 0.48 |
NGC 2232 | -258.61 | -177.32 | -41.08 | -20.31 | -12.89 | -10.70 | 0.83 | 0.61 | 0.21 | 0.49 | 0.33 | 0.08 |
OCSN 50 | -175.67 | 22.91 | -71.36 | -14.80 | -5.66 | -5.85 | 2.04 | 1.63 | 0.53 | 0.95 | 0.18 | 0.44 |
OCSN 56 | -345.38 | -124.03 | -97.58 | -26.34 | -8.58 | -8.71 | 3.05 | 1.12 | 0.95 | 1.55 | 0.61 | 0.38 |
OCSN 61 | -344.72 | -159.75 | -115.31 | -25.59 | -11.69 | -10.95 | 1.08 | 0.56 | 0.41 | 0.41 | 0.18 | 0.13 |
OCSN 64 | -268.51 | -132.13 | -122.54 | -26.17 | -5.31 | -5.58 | 11.81 | 5.84 | 5.39 | 1.31 | 0.78 | 0.68 |
Theia 69 | -115.07 | -120.16 | -21.54 | -12.56 | -18.83 | -4.54 | 0.52 | 0.77 | 0.46 | 0.81 | 0.84 | 0.22 |
Theia 71 | -177.34 | 7.91 | -57.94 | -14.92 | -6.11 | -6.07 | 0.79 | 0.85 | 0.48 | 1.77 | 0.01 | 0.44 |
Theia 72 | -212.29 | -187.58 | -26.09 | -18.24 | -22.33 | -6.41 | 1.83 | 1.40 | 0.43 | 1.13 | 0.90 | 0.20 |
Name | Other name | Age | |||||||||
[pc] | % | ||||||||||
ASCC 20 | - | 525 | 110 | 306 | [] | ||||||
ASCC 24 | - | 18 | 4 | 24 | [] | ||||||
CWNU 1057 | - | 54 | 14 | 25 | [] | ||||||
CWNU 1111 | - | 40 | 8 | 25 | [] | ||||||
HSC 1340 | - | 194 | 65 | 94 | [] | ||||||
HSC 1373 | - | 46 | 10 | 19 | [] | ||||||
HSC 1469 | - | 41 | 15 | 23 | [] | ||||||
HSC 1523 | - | 17 | 4 | 13 | [] | ||||||
HSC 1640 | Eridanus-North | 128 | 24 | 63 | [] | ||||||
HSC 1687 | - | 19 | 6 | 14 | [] | ||||||
HSC 1692 | - | 24 | 4 | 13 | [] | ||||||
NGC 2232 | - | 286 | 101 | 171 | [] | ||||||
OCSN 50 | - | 24 | 9 | 14 | [] | ||||||
OCSN 56 | omega-Ori | 52 | 5 | 23 | [] | ||||||
OCSN 61 | OBP-b | 530 | 102 | 312 | [] | ||||||
OCSN 64 | OBP-e | 67 | 11 | 48 | [] | ||||||
Theia 69 | - | 31 | 8 | 17 | [] | ||||||
Theia 71 | - | 13 | 3 | 6 | [] | ||||||
Theia 72 | MON-Theia72 | 140 | 14 | 70 | [] |
Name | 100 pc | 95 pc | 90 pc | 85 pc | 80 pc | 75 pc | 70 pc | 65 pc | 60 pc | 55 pc | 50 pc | 45 pc | 40 pc | 35 pc | 30 pc |
All clusters | 68.0 | 61.5 | 54.3 | 47.3 | 40.3 | 33.2 | 28.0 | 24.7 | 21.2 | 17.8 | 13.9 | 10.0 | 5.4 | 0.8 | 0.0 |
ASCC 20 | 23.0 | 23.0 | 23.0 | 23.0 | 23.0 | 22.9 | 22.8 | 22.0 | 19.6 | 16.9 | 13.5 | 9.8 | 5.3 | 0.8 | 0.0 |
ASCC 24 | 1.9 | 1.9 | 1.8 | 1.0 | 0.1 | - | - | - | - | - | - | - | - | - | - |
CWNU 1057 | 3.0 | 3.0 | 2.9 | 2.8 | 2.6 | 2.4 | 1.9 | 1.4 | 0.9 | 0.5 | 0.2 | 0.1 | 0.0 | 0.0 | 0.0 |
CWNU 1111 | 1.5 | 1.2 | 0.7 | 0.3 | 0.1 | 0.0 | 0.0 | - | - | - | - | - | - | - | - |
HSC 1340 | 1.1 | 0.1 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | - | - | - | - | - |
HSC 1373 | 2.4 | 1.9 | 1.1 | 0.4 | 0.1 | 0.0 | 0.0 | 0.0 | - | - | - | - | - | - | - |
HSC 1469 | 2.6 | 1.9 | 1.2 | 0.5 | 0.2 | 0.0 | 0.0 | - | - | - | - | - | - | - | - |
HSC 1523 | 1.2 | 1.0 | 0.8 | 0.6 | 0.4 | 0.3 | 0.2 | 0.1 | 0.1 | 0.0 | 0.0 | - | - | - | - |
HSC 1640 | 1.6 | 0.7 | 0.3 | 0.1 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | - | - | - | - | - | - |
HSC 1687 | 1.2 | 0.4 | 0.0 | - | - | - | - | - | - | - | - | - | - | - | - |
HSC 1692 | 1.5 | 1.4 | 1.2 | 0.7 | 0.3 | 0.1 | 0.0 | - | - | - | - | - | - | - | - |
NGC 2232 | 9.7 | 5.5 | 0.7 | - | - | - | - | - | - | - | - | - | - | - | - |
OCSN 50 | 3.2 | 3.2 | 2.8 | 1.4 | 0.2 | 0.0 | - | - | - | - | - | - | - | - | - |
OCSN 56 | 3.7 | 3.4 | 3.1 | 2.7 | 2.4 | 1.8 | 1.4 | 1.0 | 0.7 | 0.5 | 0.2 | 0.1 | 0.1 | 0.0 | 0.0 |
OCSN 61 | 29.2 | 25.9 | 21.8 | 17.5 | 12.2 | 6.3 | 1.9 | 0.3 | 0.0 | 0.0 | 0.0 | - | - | - | - |
OCSN 64 | 3.9 | 3.6 | 3.1 | 2.5 | 1.7 | 1.0 | 0.5 | 0.3 | 0.1 | 0.0 | 0.0 | 0.0 | - | - | - |
Theia 69 | 1.3 | 1.0 | 0.7 | 0.4 | 0.2 | 0.1 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
Theia 71 | 2.1 | 2.1 | 1.9 | 1.8 | 1.1 | 0.4 | 0.0 | 0.0 | 0.0 | - | - | - | - | - | - |
Theia 72 | 4.4 | 4.1 | 3.7 | 3.0 | 2.3 | 1.5 | 0.8 | 0.3 | 0.1 | 0.0 | 0.0 | 0.0 | - | - | - |
Appendix C Initial condition test
We tested the robustness of our result by varying the initial solar parameters (i.e. the Sun’s height above the disk, Galactocentric distance, and velocity respect to the Local Standard of Rest; LSR), as there is no unique definition for these quantities. We then integrated the orbits of the Solar System and the clusters and performed the same analysis described in Sect. 3.
For this test, we assumed a Galactocentric distance of (GRAVITY Collaboration et al. 2018), a vertical height of (Bennett & Bovy 2019), and a solar velocity relative to the LSR of () = (11.1, 12.24, 7.25) (Schönrich et al. 2010). We find that the past relative distances of the Sun and the clusters, and thus the resulting probabilities of having a SN within the 10Be anomaly window, remain largely unchanged over the past 20 Myr. This can be explained by the relatively short integration time and by the fact that we consider the relative distances between the clusters and the Sun rather than their absolute positions. This is also in agreement with previous studies that used similar integration periods (see e.g., Miret-Roig et al. 2020), further supporting the robustness of our conclusions.
Appendix D Systematic error study
In this section, we assess the impact of the systematic uncertainties in the geological age dating on our results. A shift in the timing of the 10Be anomaly would alter the interval during which the Solar System is close to specific clusters, potentially affecting the SN probability estimates. To test the impact of this on our results, we adopted a systematic uncertainty on the geological age dating of . This value is the largest systematic uncertainty we were able to find for the geological samples used in the paper by Koll et al. (2025) and is sourced from the work by Wallner et al. (2021) on Crust-3, one of the samples also used in Koll et al. (2025). We then repeated the SN probability analysis (see Sect. 3.2) using two shifted time windows for the 10Be anomaly: [12.0, 10.6] Myr and [11.0, 9.6] Myr.
We find that the total SN probability is only weakly affected by the assumed systematic age uncertainty. As shown in the left panel of Fig. 2, the total SN probability slightly increases for the older time window and slightly decreases for the younger one, primarily due to changes in the proximity of ASCC 20 and OCSN 61. The two main contributors are still ASCC 20 and OCSN 61. The right panel of Fig. 2 shows how the time shift affects the SN probability for these two clusters. These results confirm that our main conclusion remains robust: a nearby SN is a viable explanation for the 10Be anomaly.